5) For the following function, C computes the cost in millions of dollars of implementing a city recycling project when x percent of the citizens participate. C(x)=1.5x/100-x

a) Using this model, approximately what is the cost if 40% of the citizens participate?

b) Using this model, estimate the percentage of participation that can be excepted if $5 million is spent on this recycling project? Set up an equation and solve algebraically. Round to the nearest whole percent.


Answer by checkley71(8403)

C(X)=1.5X/(100-X)
C(40)=1.5*40/(100-40)
C(40)=60/60
C(40)=1 MILLION IS THE COST.
5=1.5X/(100-X)
5*(100-X)=1.5X
500-5X=1.5X
500=1.5X+5X
6.5X=500
X=500/6.5
X=77 % NEEDED TOP PARTICIPATE FOR THE $5 MILLION COST.